OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of –3, 5 of multiplicity 2.

OpenStudy (anonymous):

\[(x+3)^2(x-5)^2\] Expand if required.

OpenStudy (anonymous):

Is that (x+3)^2 times (x-5)^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It expands to \[x^4-4\cdot x^3-26\cdot x^2+60\cdot x+225\]

OpenStudy (anonymous):

you mean like: x^4-4x^3-26x^2+60x+225

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are they both multiplicity of two?

OpenStudy (anonymous):

if not then that solution is not correct

OpenStudy (anonymous):

Uh the way its phrased it suggests that both are of multiplicity 2.

OpenStudy (anonymous):

I took it as only being the 5, bit confusing, I think.

OpenStudy (anonymous):

Fine, if only root 5 is of multiplicity 2 then: \[(x + 3)\cdot(x - 5)^2\] Expand as necessary.

OpenStudy (anonymous):

You get: \[x^3-7x^2-5x+75\]